### v1

@@ B.2 Joint-Time-Scale Mean Square Convergence
The analysis of the slow iterate’s covariance $\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}$ involves an error term that depends on the deviation of cross covariance $C_{t}=\mathbb{E}[\tilde{x}_{t}\hat{y}_{t}^{T}]$ from $\gamma_{t}\Sigma_{fs}$. Recall the recursions in Eq. (40) derived in Section A.1, restated for reference below:
@@
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}$ (52)
@@
$\lVert R_{t}^{fs}(C_{t})\rVert\leq M_{fs}^{\prime}\left(\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert+\alpha_{t}\lVert X_{t}\rVert+\gamma_{t}\lVert\Sigma_{fs}\rVert\right)$
@@
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{fs}^{\prime}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert+\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert$
$+\gamma_{t}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert\lVert A_{sf}\rVert+\alpha_{t}\gamma_{t}M_{fs}\lVert X_{t}\rVert.$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert+\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert+\gamma_{t}^{2}M_{f}^{\prime\prime}\lVert A_{sf}\rVert+\alpha_{t}^{2}\gamma_{t}M_{f}^{\prime\prime}\lVert\Sigma_{ff}\rVert$
for some constant $M_{f}^{\prime\prime}$. Since $b<1$ and $\alpha_{t}^{2}\leq\gamma_{t}$, $\gamma_{t}^{2}$ dominates and the above is a recursion of the form
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert+M_{fs}^{\prime}\gamma_{t}^{2}$ (54)
@@
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert+M_{fs}^{\prime}\sum_{t=1}^{n}\gamma_{t}^{2}\prod_{j=t+1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}.$
@@ B.3 Slow-Time-Scale Mean Square Convergence
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
@@
$\lVert R_{t}^{fs}(Y_{t})\rVert\leq M_{s}^{\prime}\left(\lVert Y_{t}\rVert+\alpha_{t}\lVert C_{t}\rVert+\alpha_{t}\lVert X_{t}\rVert\right).$
@@
$\lVert R_{t}^{fs}(Y_{t})\rVert\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}\rVert+\alpha_{t}^{2}\right)\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert+\gamma_{t}\lVert\Sigma_{ss}\rVert+\alpha_{t}^{2}\right)$
@@
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert$ $\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{2}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert+\frac{\gamma_{t}}{t}\lVert\Sigma_{ss}\rVert$
$+\gamma_{t}\left\lVert(C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right\rVert+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\left(\gamma_{t}\lVert\Sigma_{ss}\rVert+\alpha_{t}^{2}\right).$
@@
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert+\frac{\gamma_{t}}{t}M_{s}^{\prime\prime},$ (58)
@@
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
@@ Appendix C Non-Asymptotic CLT for TSA-PR
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+n^{a-b/2}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+n^{a-b/2}+n^{b/2}\right)\right),\end{split}$ (60)
@@
We have from Theorem 3.1 and Jensen’s inequality that
$\mathbb{E}\lVert x_{t}\rVert=\mathcal{O}(\sqrt{\alpha_{t}}+\sqrt{\gamma_{t}}),\quad\mathbb{E}\lVert y_{t}\rVert=\mathcal{O}(\sqrt{\gamma_{t}}).$
Substituting, we obtain
$\mathbb{E}\lVert\sum_{t=1}^{n}\alpha_{t}^{-1}(x_{t}-x_{t+1})\rVert+\mathbb{E}\lVert\sum_{t=1}^{n}\gamma_{t}^{-1}(y_{t}-y_{t+1})\rVert=\mathcal{O}\left(n^{a/2}+n^{a-b/2}+n^{b/2}\right).$
Combining with Lemma 5.1 and 5.2 yields Theorem 3.2. Similar to how we included the transient term $\sqrt{\gamma_{t}}$ for $\mathbb{E}\lVert x_{t}\rVert$ to obtain the $n^{a-b/2}$ in the above equation, a refined inspection of all rates in the proof of Theorem 3.1 illustrate how to control the transient terms in the Wasserstein-1 distance for the Polyak-Ruppert averaged errors in Theorem 3.2.

### v2

@@ B.2 Joint-Time-Scale Mean Square Convergence
The analysis of the slow iterate’s covariance $\mathbb{E}\hat{y}_{t}\hat{y}_{t}^{T}$ involves an error term that depends on the deviation of cross covariance $C_{t}=\mathbb{E}[\tilde{x}_{t}\hat{y}_{t}^{T}]$ from $\gamma_{t}\Sigma_{fs}$. Because $x,y$ can have different dimensions, the contraction is proved for the cross covariance in the norm
$\lVert C\rVert_{P_{\Delta},P_{ff}}^{2}=\sup_{v\neq 0}\frac{v^{T}C^{T}P_{ff}Cv}{v^{T}P_{\Delta}v}.$
Recall the recursions in Eq. (41) derived in Section A.1, restated for reference below:
@@
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert\leq\sqrt{\kappa(P_{ff})\kappa(P_{\Delta})}\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}$ (54)
@@
$\lVert R_{t}^{fs}(C_{t})\rVert_{P_{\Delta},P_{ff}}\leq M_{fs}^{\prime}\left(\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+\alpha_{t}\lVert X_{t}\rVert+\gamma_{t}\lVert\Sigma_{fs}\rVert\right)$
@@
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{2}+\gamma_{t}M_{fs}^{\prime}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$
$+\gamma_{t}\lVert X_{t}-\alpha_{t}\Sigma_{ff}\rVert_{P_{ff}}\lVert A_{sf}\rVert_{P_{\Delta},P_{ff}}+\gamma_{t}M_{fs}\left(\alpha_{t}\lVert X_{t}\rVert+\gamma_{t}\lVert\Sigma_{fs}\rVert\right).$
$\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime\prime}\left(\frac{\gamma_{t}}{t}\lVert\Sigma_{fs}\rVert+\gamma_{t}^{2}\lVert A_{sf}\rVert+\alpha_{t}^{2}\gamma_{t}\lVert\Sigma_{ff}\rVert\right)$
for some constant $M_{fs}^{\prime\prime}$. Since $b<1$ and $\alpha_{t}^{2}\leq\gamma_{t}$, $\gamma_{t}^{2}$ dominates and the above is a recursion of the form
$\lVert C_{t+1}-\gamma_{t+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}\leq\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{t}-\gamma_{t}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime}\gamma_{t}^{2}$ (56)
@@
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}^{\prime}\sum_{t=1}^{n}\gamma_{t}^{2}\prod_{j=t+1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)$
$\leq\prod_{t=1}^{n}\left(1-\alpha_{t}\frac{\mu_{ff}}{4}\right)\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}+M_{fs}\frac{\gamma_{n}^{2}}{\alpha_{n}}.$
The unweighted norm is recovered by the equivalence of norms
$\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert$ $\leq\sqrt{\frac{\lambda_{\max}(P_{ff}^{-1})}{\lambda_{\min}(P_{\Delta}^{-1})}}\lVert C_{n+1}-\gamma_{n+1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$
$\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert_{P_{\Delta},P_{ff}}$ $\leq\sqrt{\frac{\lambda_{\max}(P_{ff})}{\lambda_{\min}(P_{\Delta})}}\lVert C_{1}-\gamma_{1}\Sigma_{fs}\rVert.$
@@ B.3 Slow-Time-Scale Mean Square Convergence
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
@@
$\lVert R_{t}^{ss}(Y_{t})\rVert_{P_{\Delta}}\leq M_{s}^{\prime}\left(\lVert Y_{t}\rVert_{P_{\Delta}}+\alpha_{t}\lVert A_{sf}C_{t}\rVert_{P_{\Delta}}+\alpha_{t}\lVert X_{t}\rVert_{P_{\Delta}}\right).$
@@
$\lVert R_{t}^{ss}(Y_{t})\rVert_{P_{\Delta}}\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right)\leq M_{s}^{\prime\prime}\left(\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\gamma_{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right)$
@@
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert_{P_{\Delta}}$ $\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{2}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\frac{\gamma_{t}}{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}$
$+\gamma_{t}\left\lVert(C_{t}-\gamma_{t}\Sigma_{fs})^{T}A_{sf}^{T}+A_{sf}(C_{t}-\gamma_{t}\Sigma_{fs})\right\rVert_{P_{\Delta}}+\frac{\gamma_{t}^{2}}{\alpha_{t}}M_{s}^{\prime}\left(\gamma_{t}\lVert\Sigma_{ss}\rVert_{P_{\Delta}}+\alpha_{t}^{2}\right).$
@@
$\lVert Y_{t+1}-\gamma_{t+1}\Sigma_{ss}\rVert_{P_{\Delta}}\leq\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{t}-\gamma_{t}\Sigma_{ss}\rVert_{P_{\Delta}}+\frac{\gamma_{t}}{t}M_{s}^{\prime\prime},$ (60)
@@
$\lVert Y_{n+1}-\gamma_{n+1}\Sigma_{ss}\rVert\leq\kappa(P_{\Delta})\prod_{t=1}^{n}\left(1-\gamma_{t}\frac{\mu_{\Delta}}{4}\right)\lVert Y_{1}-\gamma_{1}\Sigma_{ss}\rVert+\frac{M_{s}}{n}.$
@@ Appendix C Non-Asymptotic CLT for TSA-PR
$\begin{split}d_{1}\left(\sqrt{n}G(\bar{x}_{n}-x^{*}),(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\\ d_{1}\left(\sqrt{n}\Delta(\bar{y}_{n}-y^{*}),(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/2}Z_{2}\right)&=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right),\end{split}$ (62)
@@
Equivalently,
$\lim_{n\to\infty}\mathrm{Cov}G\bar{x}_{n}=\mathrm{Cov}(W_{1}-A_{fs}A_{ss}^{-1}V_{1}),\quad\lim_{n\to\infty}\mathrm{Cov}\Delta\bar{y}_{n}=\mathrm{Cov}(V_{1}-A_{sf}A_{ff}^{-1}W_{1}).$ (64)
@@
Next, we obtain bounds on $\mathbb{E}\lVert x_{t}\rVert$ and $\mathbb{E}\lVert y_{t}\rVert$ from Theorem 3.1 and Jensen’s inequality. Following the proof of Theorem 3.1, the dominant rates that appear in the error bounds for TSA (without PR averaging) are given by
$\lVert\mathbb{E}(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})^{T}-\alpha_{t+1}\Sigma_{ff}\rVert=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}\right)$
$\lVert\mathbb{E}(x_{t+1}-x_{\infty}(y_{t+1})+L_{t}\hat{y}_{t+1})\hat{y}_{t+1}^{T}-\gamma_{t+1}\Sigma_{fs}\rVert=\mathcal{O}\left(\alpha_{t}^{-1}\left(\frac{\gamma_{t}}{t}+\gamma_{t}^{2}+\alpha_{t}^{2}\gamma_{t}\right)+\frac{\gamma_{t}}{t}+\gamma_{t}^{2}\right)$
$\lVert\mathbb{E}\hat{y}_{t+1}\hat{y}_{t+1}^{T}-\gamma_{t+1}\Sigma_{ss}\rVert=\mathcal{O}\left(\frac{1}{t}+\alpha_{t}\gamma_{t}\right).$
We are interested in obtaining $\mathbb{E}\lVert x_{t}\rVert^{2}$ and $\mathbb{E}\lVert y_{t}\rVert^{2}$ when $x^{*},y^{*}$ are zero (due to zero bias assumed for simplicity). Hence, $\hat{y}_{t}=y_{t}$ and we have the square norm of the iterates
$\lVert\mathbb{E}y_{t+1}y_{t+1}^{T}\rVert$ $=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right)$
$\lVert\mathbb{E}x_{t+1}x_{t+1}^{T}\rVert$ $=\mathcal{O}\left(\alpha_{t}+\gamma_{t}+\frac{1}{t}+\lVert\mathbb{E}x_{t+1}y_{t+1}\rVert+\mathbb{E}\lVert y_{t+1}y_{t+1}^{T}\rVert\right).$
Combining, it holds that
$\mathrm{Tr}\mathbb{E}x_{t+1}x_{t+1}^{T}$ $\leq\mathrm{dim}(x)\lVert\mathbb{E}x_{t+1}x_{t+1}^{T}\rVert=\mathcal{O}\left(\alpha_{t}+\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right),$
$\mathrm{Tr}\mathbb{E}y_{t+1}y_{t+1}^{T}$ $\leq\mathrm{dim}(y)\mathbb{E}\lVert y_{t+1}y_{t+1}^{T}\rVert=\mathcal{O}\left(\gamma_{t}+\frac{1}{t}+\alpha_{t}\gamma_{t}\right)$
excluding some non-dominant terms when $b=a+1/\log n$ and $a=1/2+1/\log n$ for large $n$. Using $\mathbb{E}\lVert x_{t}\rVert\leq\sqrt{\mathrm{Tr}\mathbb{E}x_{t}x_{t}^{T}}$, we obtain bounds on $\mathbb{E}\lVert x_{t}\rVert$ and $\mathbb{E}\lVert y_{t}\rVert$.
To complete the error bound, we substitute the above into Eqs. (70) and (71) and simplify as
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}t^{a-1}\mathbb{E}\lVert x_{t}\rVert=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{a/2}+\frac{n^{a-1/2}}{a-1/2}+\frac{1}{b-a}\right)\right),$
$\frac{1}{\sqrt{n}}\sum_{t=1}^{n}t^{b-1}\mathbb{E}\lVert y_{t}\rVert=\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(n^{b/2}+\frac{n^{b-1/2}}{b-1/2}+\frac{n^{(b-a)/2}}{b-a}\right)\right).$
Combining, $d_{1}\left(\sqrt{n}G\bar{x}_{n},(G\bar{\Sigma}_{ff}G^{T})^{1/2}Z_{1}\right)$ and $d_{2}\left(\sqrt{n}(\Delta\bar{y}_{n},(\Delta\bar{\Sigma}_{ss}\Delta^{T})^{1/}Z_{2}\right)$ are dominated by the sum of the above terms which is simplified as
$\mathcal{O}\left(\frac{1}{\sqrt{n}}\left(\frac{n^{a-1/2}}{a-1/2}+\frac{n^{(b-a)/2}}{b-a}+n^{b/2}\right)\right).$
Therefore, with the choice of step sizes $a=1/2+1/\log n$ and $b=a+1/\log n$ when $n$ is sufficiently large, we obtain the error bound $\mathcal{O}(n^{-1/4})$.